NvidiaIntely (nvidiaintely):

A compound was analyzed and found to contain only 14 grams of calcium, 11 grams of oxygen and roughly 0.7 grams of hydrogen. What is the empirical formula of the compound? Ca2OH CaOH CaO2H Ca(OH)2

NvidiaIntely (nvidiaintely):

@welshfella @Kevin @Will.H @brainlesslonelygirl @brainzonly (DON"T COME BRAINZONLY) @Photon336

NvidiaIntely (nvidiaintely):

Can someone reply please?

OpenStudy (brainlesslonelygirl):

last one

NvidiaIntely (nvidiaintely):

What do you mean?

NvidiaIntely (nvidiaintely):

Well if someone could explain to me in very much detail, I would love that. But I cannot find anyone who can right now. @welshfella, can you help me?

OpenStudy (jiteshmeghwal9):

OK nvm those replies

NvidiaIntely (nvidiaintely):

What replies?

OpenStudy (welshfella):

divide each mass by the relative atomic mass of the element Ca :- 14 / 40 O :- 11 / 16 H:- 0.7 / 1.008

OpenStudy (brainlesslonelygirl):

for the empirical formula you just simply take the given mass of the element and divide it by the atomic mass of that particular element. for example in this question, for calcium 14 (which is the given mass)/40 (which is the atomic mass of calcium, you can find this in the periodic table)= 0.35 repeat this for all the elements. oxygen= 11/16= 0.68 hydrogen= 0.7/1 = 0.7 now we see which answer is the smallest and we divide the rest by that. in this case the smallest answer is of calcium, 0.35 so we divide all three answers by 0.35. and we get calcium= 0.35/0.35=1 oxygen=0.68/0.35= 1.9 which can be rounded to give 2 hydrogen= 0.7/0.35= 2 now we have empirical formula, which is he proportions of the elements in the given compounds. so now we know the proportion for calcium is 1, O is 2 and H is 2 aswell now we can get the formula for the compound which is Ca(OH)2

NvidiaIntely (nvidiaintely):

Thanks @brainlesslonelygirl! Your great at explaining.

OpenStudy (welshfella):

isn't Ca = 40??

NvidiaIntely (nvidiaintely):

idk

OpenStudy (brainlesslonelygirl):

yea it is 40. why?

OpenStudy (welshfella):

yes to be more exact its 40.08 not 35

OpenStudy (brainlesslonelygirl):

and you're welcome @NvidiaIntely glad to be of help

OpenStudy (welshfella):

Oh sorry - I see now that youve used 40...

OpenStudy (brainzonly):

Why tag me if you don't want me to come?